OpenStudy (anonymous):

Solve. 2x2 – 8x – 14 = 0 A.x = 4 ± 3√2 B.x = 2 ± 2√7 C.x = 4 ± √3 D.x = 2 ± √11

OpenStudy (anonymous):

Use quadratic equation to solve for x.

OpenStudy (phi):

Did you try? the first thing I would do is factor out a 2 from each term 2(x^2-4x-7)=0 divide both sides by 2 x^2 -4x-7= 0/2 or x^2 -4x-7= 0 now do the quadratic formula See http://www.khanacademy.org/math/algebra/quadratics/v/applying-the-quadratic-formula